I have to do the following problem and I'm not sure if I have it correct.
Is -4 a solution to the equation 7x - 5 + 3x = 6 + x - 10?
This is what I have done:7(-4) - 5 + 3 (-4) = 6 + (-4) - 10
28 - 5 + -12 = 10 - 10
23 + -12 = 10 - 10
-11 = 0
So my answer is No it is not a solution to the equation. Is this correct? Please Help!!!!

Answer by jake_6233(74)

7x - 5 + 3x = 6 + x - 10? Combining like terms on both sides
10x-5=x-4
9x=1
x=1/9 to check your answer
7(1/9)-5 +3(1/9)= 6+(1/9) -10
7/9 + 3/9 -45/9 = 54/9 +1/9 -90/9
(-35/9) = (-35/9)
I hope that this helps you. Please let me know if I can assist you further.